# Reactive capacitance of flat patches of arbitrary shape

### v1

@@ I Introduction
Figure 1: (a) A domain $\Omega$ with reflecting boundary $\partial\Omega_{0}$ (in gray) A flat patch $\Gamma$ on the horizontal plane of the upper half-space.
@@ II.1 Exterior Steklov problem
The first eigenvalue $\mu_{0}$ is simple and strictly positive, whereas the associated eigenfunction $\Psi_{0}$ does not change sign. The restrictions of the Steklov eigenfunctions onto the patch $\Gamma$, $\Psi_{k}|_{\Gamma}$, form a complete orthonormal basis of the $L^{2}(\Gamma)$ space of square-integrable functions on $\Gamma$:
$\int\limits_{\Gamma}\Psi_{k}(\bm{y})\,\Psi_{k^{\prime}}(\bm{y})d\bm{y}=\delta_{k,k^{\prime}}.$ (5)
(see [74, 75, 76] for a broad overview of spectral properties of the conventional interior Steklov problem).
@@
where $\bar{\bm{y}}^{\prime}$ is the mirror reflection of $\bm{y}^{\prime}$ with respect to the horizontal plane. Multiplying Eq. (3a) by $G(\bm{y},\bm{y}^{\prime})$, multiplying Eq. (6a) by $\Psi_{k}(\bm{y})$, subtracting these equations, integrating them over $\bm{y}\in{\mathbb{R}}^{3}_{+}$, using the Green’s formula and the boundary conditions, one finds
@@
We emphasize that the domain monotonicity property does not hold for the Steklov problem in a more general setting of the exterior of a compact set. For instance, if $\Gamma$ is an oblate spheroidal surface with semi-axes $a,a,b$, the domain monotonicity property (13) fails, as illustrated in Fig. 7 of [80], which shows the non-monotonous dependence of the principal Steklov eigenvalue on the aspect ratio $a/b$. This example highlights that the domain monotonicity property for patches strongly relies on the flatness of these domains.
@@ II.2 Spectral expansion and basic properties
(here we employ the convention that the capacitance of a ball of radius $R$ is equal to $R$).
@@ II.4 Sigmoidal approximation and several bounds
we retrieve the sigmoidal approximation for the reactive capacitance. This relation was proposed in [68] as an empirical approximation that correctly reproduces both small-$\mu$ and large-$\mu$ limits of $C(\mu)$. In turn, it follows here directly from the spectral expansion (25) by neglecting its last term, thus allowing one to control its accuracy (see below). Note that a similar approximation appeared in [90] in the analysis of the principal eigenvalue of the Laplace operator with Robin boundary condition, as well as in [37, 65]. In the special case of a circular patch, the approximation (27) was shown to be accurate over the entire range of $\mu$ within the maximal relative error of $4\%$ [68]. One of the aims of the present work consists in extending this result to patches of arbitrary shape. For this purpose, we need to compute numerically the Steklov eigenvalues $\mu_{k}$ and the spectral weights $F_{k}$ (see Sec. III).
@@
Taking the limit $\mu\to 0$ and using the Taylor expansion (17) with two terms, $C(\mu)\approx c_{1}\mu-c_{2}\mu^{2}+O(\mu^{3})$, one can compute the third term in the right-hand side of Eq. (28) as
@@ II.5 Probabilistic interpretations
At each arrival onto the patch $\Gamma$, the diffusing particle modeled by reflected Brownian motion may react and thus disappear. For a constant reactivity parameter $\mu$, the reaction event is triggered when the number of arrivals (i.e., failed reaction attempts), represented by $\ell_{t}$, exceeds a random threshold $\hat{\ell}$ that obeys the exponential distribution: ${\mathbb{P}}\{\hat{\ell}>\ell\}=e^{-\mu\ell}$ (see [64] for details). In this way, the first-reaction time (FRT) on the patch $\Gamma$ is defined as the first-crossing time of the threshold $\hat{\ell}$ by $\ell_{t}$: $\tau=\inf\{t>0~:~\ell_{t}<\hat{\ell}\}$. Since the reflected Brownian motion in ${\mathbb{R}}^{3}_{+}$ is transient, there is a finite probability of escaping to infinity without reacting on the patch; in this case, the FRT is infinite: $\tau=+\infty$. In turn, the probability of the reaction event, ${\mathbb{P}}_{\bm{y}}\{\tau<+\infty\}$, is known to satisfy Eq. (1) and is thus given by $w(\bm{y};\mu)$. We conclude that the auxiliary function $w(\bm{y};\mu)$ admits a simple interpretation as the probability of reaction on the patch $\Gamma$ with the reactivity parameter $\mu$, when the starting point is $\bm{y}$. In turn, $1-w(\bm{y};\mu)$ is the escape probability without reaction. We also note that the reaction time $\tau$ is finite if and only if the limit $\ell_{\infty}$ is above the threshold $\hat{\ell}$, i.e., the reaction event $\{\tau<+\infty\}$ is identical to $\{\ell_{\infty}>\hat{\ell}\}$. As a consequence, we get
@@
$=\int\limits_{0}^{\infty}d\mu\,\underbrace{\mu e^{-\mu\ell}}_{\textrm{PDE of}~\hat{\ell}}~\cdot~\underbrace{{\mathbb{P}}_{\bm{y}}\{\ell_{\infty}>\ell\}}_{\int\nolimits_{\ell}^{\infty}\rho_{\infty}(\ell^{\prime}|\bm{y})d\ell^{\prime}}$
$=1-\int\limits_{0}^{\infty}d\mu\,e^{-\mu\ell}\,\rho_{\infty}(\ell|\bm{y})=1-{\mathbb{E}}_{\bm{y}}\{e^{-\mu\ell_{\infty}}\},$
@@
$P_{\circ}(\mu)=\int\limits_{\Gamma}(1-w(\bm{y};\mu))\frac{d\bm{y}}{|\Gamma|}=\frac{2\pi C(\mu)}{\mu|\Gamma|}\,$, (42)
where we used the boundary condition (1b), the divergence theorem, and the asymptotic behavior (1d) at infinity (the subscript $\circ$ refers to the uniform distribution of the starting point on $\Gamma$). Inverting this relation, one can express the reactive capacitance $C(\mu)$ in terms of the escape probability on the patch. In turn, using Eq. (41), we can relate $C(\mu)$ to the moment-generating function:
${\mathbb{E}}_{\circ}\{e^{-\mu\ell_{\infty}}\}=\frac{2\pi C(\mu)}{\mu|\Gamma|}\,$, (43)
@@
${\mathbb{E}}_{\circ}\{e^{-\mu\ell_{\infty}}\}=\int\limits_{\Gamma}{\mathbb{E}}_{\bm{y}}\{e^{-\mu\ell_{\infty}}\}\,\frac{d\bm{y}}{|\Gamma|}$ (44)
@@
${\mathbb{E}}_{\circ}\{\ell_{\infty}^{n}\}=\frac{2\pi}{|\Gamma|}c_{n+1}\quad(n=0,1,\ldots).$ (45)
For instance, we retrieve $c_{1}=|\Gamma|/(2\pi)$ at $n=0$, whereas $c_{2}$ is proportional to the mean boundary local time. Moreover, substituting the spectral expansion (14) into Eq. (43) and performing the inverse Laplace transform with respect to $\mu$, one gets the probability density of $\ell_{\infty}$:
$\rho_{\infty}(\ell|\circ)=\sum\limits_{k=0}^{\infty}\mu_{k}F_{k}e^{-\mu_{k}\ell}$ (46)
@@ II.6 Domain monotonicity
It is intuitively appealing that the same statement should hold for partially reactive patches, i.e., a larger patch captures diffusing particles more efficiently. In this subsection, we establish this domain monotonicity property and some other monotonicity relations.
@@
$\int\limits_{\Gamma}\partial_{n}\tilde{w}=0,$ (65)
@@
$0$ $=\int\limits_{{\mathbb{R}}^{3}_{+}}\tilde{w}\,\Delta\tilde{w}=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}+\int\limits_{\Gamma}\tilde{w}\,\partial_{n}\tilde{w}$.
@@
$0$ $=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}-\mu\int\limits_{\Gamma_{i}}\tilde{w}^{2}+\biggl(\frac{\mu}{C(\mu)}-\frac{\mu}{C(\infty)}\biggr)\int\limits_{\Gamma}\tilde{w}$
$-2\pi\int\limits_{\Gamma}\tilde{w}\,h\,.$
@@
$\int\limits_{\Gamma}\tilde{w}\,h=\int\limits_{\Gamma}\underbrace{w(\bm{y};\infty)}_{=1}\partial_{n}\tilde{w}=0,$ (67)
@@
$=\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}+\mu\int\limits_{\Gamma}\tilde{w}^{2}.$ (68)
@@
$\frac{\eta_{1}(\eta_{1}-2\pi/\mu)}{|\Gamma_{1}|}\leq\frac{\eta_{2}(\eta_{2}-2\pi/\mu)}{|\Gamma_{2}|}$ (70)
@@ II.7 Large-reactivity limit
$\mu\biggl(\frac{1}{C(\mu)}-\frac{1}{C(\infty)}\biggr)=2\pi\sum\limits_{k=0}^{\infty}\frac{\mu[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}+\mu}\propto\ln(\mu)\to\infty$ (80)
@@
$\times\underbrace{\sum\limits_{k=0}^{\infty}\Psi_{k}^{N}(\bm{y}_{1})\Psi_{k}^{N}(\bm{y}_{2})}_{=\delta(\bm{y}_{1}-\bm{y}_{2})}=\int\limits_{\Gamma}d\bm{y}\,h^{2}(\bm{y})=\infty.$
In other words, the squared harmonic measure density diverges logarithmically due to the edge singularity. This is exemplified by Eq. (74) that provides an explicit form of $h(\bm{y})$ for elliptic patches. In summary, even though the coefficients $\Psi_{k}^{N}(\infty)]^{2}$ determine the relative contributions of the Steklov eigenfunctions $\Psi_{k}^{N}$ to the reactive capacitance via Eq. (25), in analogy to the coefficients $F_{k}$ in Eq. (14), the latter are summed to $1$ that facilitates their interpretation as relative weights of the Steklov eigenfunctions $\Psi_{k}$.
@@ II.8 Bounds for the principal eigenvalue
$\int\limits_{\Gamma}\Psi_{0}\leq\sqrt{2\pi c_{n+1}\mu_{0}^{n}}\,,\qquad n=1,2,\ldots$ (88)
@@
where $E(x)$ is the complete elliptic integral of the second kind. Substituting the values of $C(\infty)=2/\pi$, $c_{2}$ and $c_{3}$ into Eqs. (83, 84), we get three upper bounds: $1.263$ (Payne’s bound), $1.178$ ($n=1$), and $1.170$ ($n=2$), whereas the numerical value of the principal eigenvalue is $1.159$. One sees that the bound becomes tighter as $n$ increases, as expected from the spectral form (21). Using the numerical value $F_{0}\approx 0.978$, we get the lower bounds $1.152$ and $1.157$, where the last value is very close to the numerical value of $\mu_{0}$. We recall that $c_{2}$ and $c_{3}$ for arbitrary patches are expressed via Eq. (18) in terms of the function $\omega_{\Gamma}(\bm{x})$ from Eq. (19). In Appendix B, we compute $\mathcal{A}_{\Gamma}$ for rectangular and elliptic patches; in the latter case, we also compare the upper bounds (83) and (86).
@@ III.1 Steklov eigenpairs
We complete this discussion by illustrating the difference between the Steklov problems (3) and (23). Figure 4 presents the first 9 eigenfunctions $\Psi_{k}^{N}|_{\Gamma}$ for the elliptic patch of semiaxes $1$ and $0.5$. Their comparison with panels of Fig. 3 indicates that, apart from the principal eigenfunctions $\Psi_{0}$ and $\Psi_{0}^{N}$, which are substantially different, the other eigenfunctions look quite similar, up to matching their indices. Moreover, antisymmetric eigenfunctions are actually identical, up to their signs, e.g. $\Psi_{1}^{N}=-\Psi_{1}$, $\Psi_{3}^{N}=-\Psi_{2}$, etc. This is not surprising (see Appendix A for details): as the kernels $\mathcal{G}(\bm{y},\bm{y}^{\prime})$ and $\mathcal{G}^{N}(\bm{y},\bm{y}^{\prime})$ differ by a function $\omega_{\Gamma}(\bm{y})/|\Gamma|+\omega_{\Gamma}(\bm{y}^{\prime})/|\Gamma|-\mathcal{A}_{\Gamma}$, which is symmetric on the elliptic patch, the antisymmetric eigenfunctions of these kernels should be identical.
@@
Figure 7: First 15 Steklov eigenfunctions $\Psi_{k}$ restricted to the patch formed by two disks of radii $1$ and $0.5$ (the associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the titles). The eigenfunctions were obtained by using a mesh with 964 triangles and 536 nodes.

### v2

@@ I Introduction
Figure 1: Steady-state diffusion in the upper half-space ${\mathbb{R}}^{3}_{+}$ towards a flat reactive patch $\Gamma$ on the horizontal reflecting plane.
@@ II.1 Exterior Steklov problem
The first eigenvalue $\mu_{0}$ is simple (of multiplicity $1$) and strictly positive, whereas the associated eigenfunction $\Psi_{0}$ does not change sign. The restrictions of the Steklov eigenfunctions onto the patch $\Gamma$, $\Psi_{k}|_{\Gamma}$, form a complete orthonormal basis of the $L^{2}(\Gamma)$ space of square-integrable functions on $\Gamma$:
$\int\limits_{\Gamma}\Psi_{k}(\bm{y})\,\Psi_{k^{\prime}}(\bm{y})d\bm{y}=\delta_{k,k^{\prime}}$ (5)
(see [74, 75, 76] for a broad overview of spectral properties of the conventional Steklov problem for bounded domains).
@@
where $\bar{\bm{y}}^{\prime}=(y^{\prime}_{1},y^{\prime}_{2},-y^{\prime}_{3})$ is the mirror reflection of $\bm{y}^{\prime}=(y^{\prime}_{1},y^{\prime}_{2},y^{\prime}_{3})$ with respect to the horizontal plane. Multiplying Eq. (3a) by $G(\bm{y},\bm{y}^{\prime})$, multiplying Eq. (6a) by $\Psi_{k}(\bm{y})$, subtracting these equations, integrating them over $\bm{y}\in{\mathbb{R}}^{3}_{+}$, using the Green’s formula and the boundary conditions, one finds
@@
We emphasize that the domain monotonicity property does not hold for the Steklov problem in a more general setting of the exterior of a compact set. For instance, if $\Gamma$ is an oblate spheroidal surface with semi-axes $a,a,b$, the domain monotonicity property (13) fails, as illustrated in Fig. 7 of [80], which shows the non-monotonous dependence of the principal Steklov eigenvalue on the aspect ratio $a/b$. This example highlights that the domain monotonicity property holds for patches due to their flatness.
@@ II.2 Spectral expansion and basic properties
(here we employ the convention that the electrostatic capacitance of a ball of radius $R$ is equal to $R$).
@@ II.4 Sigmoidal approximation and several bounds
we retrieve the sigmoidal approximation for the reactive capacitance. This relation was proposed in [68] as an empirical approximation that correctly reproduces both small-$\mu$ and large-$\mu$ limits of $C(\mu)$. In turn, it follows here directly from the spectral expansion (25) by neglecting its last term, thus allowing one to control its accuracy (see below). Note that a similar approximation appeared in [90] in the analysis of the principal eigenvalue of the Laplace operator with Robin boundary condition, as well as in [37, 65]. In the special case of a circular patch, the approximation (27) was shown to be accurate over the entire range of $\mu$ within the maximal relative error of $4\%$ [68]. One of the aims of the present work consists in verifying the accuracy of the sigmoidal approximation for patches of arbitrary shape. For this purpose, we need to compute numerically the Steklov eigenvalues $\mu_{k}$ and the spectral weights $F_{k}$ (see Sec. III).
@@
Taking the limit $\mu\to 0$ and using the Taylor expansion (17) with two terms, $C(\mu)\approx c_{1}\mu-c_{2}\mu^{2}+O(\mu^{3})$, one computes the third term in the right-hand side of Eq. (28) as
@@ II.5 Probabilistic interpretations
At each arrival onto the patch $\Gamma$, the diffusing particle modeled by reflected Brownian motion may react and thus disappear. For a constant reactivity parameter $\mu$, the reaction event is triggered when the number of arrivals (i.e., failed reaction attempts), represented by $\ell_{t}$, exceeds a random threshold $\hat{\ell}$ that obeys the exponential distribution: ${\mathbb{P}}\{\hat{\ell}>\ell\}=e^{-\mu\ell}$ (see [64] for details). In this way, the first-reaction time (FRT) on the patch $\Gamma$ is defined as the first-crossing time of the threshold $\hat{\ell}$ by $\ell_{t}$: $\tau=\inf\{t>0~:~\ell_{t}>\hat{\ell}\}$. Since the reflected Brownian motion in ${\mathbb{R}}^{3}_{+}$ is transient, there is a finite probability of escaping to infinity without reacting on the patch; in this case, the FRT is infinite: $\tau=+\infty$. In turn, the probability of the reaction event, ${\mathbb{P}}_{\bm{y}}\{\tau<+\infty\}$, is known to satisfy Eq. (1) and is thus given by $w(\bm{y};\mu)$. We conclude that the auxiliary function $w(\bm{y};\mu)$ admits a simple interpretation as the probability of reaction on the patch $\Gamma$ with the reactivity parameter $\mu$, when the starting point is $\bm{y}$. In turn, $1-w(\bm{y};\mu)$ is the escape probability without reaction. We also note that the reaction time $\tau$ is finite if and only if the limit $\ell_{\infty}$ is above the threshold $\hat{\ell}$, i.e., the reaction event $\{\tau<+\infty\}$ is identical to $\{\ell_{\infty}>\hat{\ell}\}$. As a consequence, we get
@@
$=\int\limits_{0}^{\infty}\underbrace{\mu e^{-\mu\ell}}_{\textrm{PDF of}~\hat{\ell}}~\cdot~\underbrace{{\mathbb{P}}_{\bm{y}}\{\ell_{\infty}>\ell\}}_{\int\nolimits_{\ell}^{\infty}\rho_{\infty}(\ell^{\prime}|\bm{y})d\ell^{\prime}}d\ell$
$=1-\int\limits_{0}^{\infty}e^{-\mu\ell}\,\rho_{\infty}(\ell|\bm{y})\,d\ell=1-{\mathbb{E}}_{\bm{y}}\{e^{-\mu\ell_{\infty}}\},$
@@
$P_{u}(\mu)=\int\limits_{\Gamma}(1-w(\bm{y};\mu))\frac{d\bm{y}}{|\Gamma|}=\frac{2\pi C(\mu)}{\mu|\Gamma|}\,$, (42)
where we used the boundary condition (1b), the divergence theorem, and the asymptotic behavior (1d) at infinity (the subscript $u$ refers to the uniform distribution of the starting point on $\Gamma$). Inverting this relation, one can express the reactive capacitance $C(\mu)$ in terms of the escape probability on the patch. In turn, using Eq. (41), we can relate $C(\mu)$ to the moment-generating function:
${\mathbb{E}}_{u}\{e^{-\mu\ell_{\infty}}\}=\frac{2\pi C(\mu)}{\mu|\Gamma|}\,$, (43)
@@
${\mathbb{E}}_{u}\{e^{-\mu\ell_{\infty}}\}=\int\limits_{\Gamma}{\mathbb{E}}_{\bm{y}}\{e^{-\mu\ell_{\infty}}\}\,\frac{d\bm{y}}{|\Gamma|}$ (44)
@@
${\mathbb{E}}_{u}\{\ell_{\infty}^{n}\}=\frac{2\pi}{|\Gamma|}c_{n+1}\quad(n=0,1,\ldots).$ (45)
For instance, we retrieve $c_{1}=|\Gamma|/(2\pi)$ at $n=0$, whereas $c_{2}$ is proportional to the mean boundary local time. As a consequence, we get
$\mathcal{A}_{\Gamma}=\frac{{\mathbb{E}}_{u}\{\ell_{\infty}\}}{|\Gamma|}\,.$ (46)
Curiously, the same constant $\mathcal{A}_{\Gamma}$ was shown to determine the asymptotic behavior of the variance of the boundary local time $\ell_{t}$ on a small patch for restricted diffusion in a bounded domain [66]. Moreover, substituting the spectral expansion (14) into Eq. (43) and performing the inverse Laplace transform with respect to $\mu$, one gets the probability density of $\ell_{\infty}$:
$\rho_{\infty}(\ell|u)=\sum\limits_{k=0}^{\infty}\mu_{k}F_{k}e^{-\mu_{k}\ell}$ (47)
@@ II.6 Domain monotonicity
It is intuitively expected that the same statement holds for partially reactive patches, i.e., a larger patch captures diffusing particles more efficiently. In this subsection, we establish this domain monotonicity property and some other monotonicity relations.
@@
$\int\limits_{\Gamma}\partial_{n}\tilde{w}\,d\bm{y}=0,$ (66)
@@
$0=\int\limits_{{\mathbb{R}}^{3}_{+}}\tilde{w}\,\Delta\tilde{w}\,d\bm{y}=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}d\bm{y}+\int\limits_{\Gamma}\tilde{w}\,\partial_{n}\tilde{w}\,d\bm{y}.$
@@
$0$ $=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}\,d\bm{y}-\mu\int\limits_{\Gamma_{i}}\tilde{w}^{2}\,d\bm{y}$
$+\biggl(\frac{\mu}{C(\mu)}-\frac{\mu}{C(\infty)}\biggr)\int\limits_{\Gamma}\tilde{w}\,d\bm{y}-2\pi\int\limits_{\Gamma}\tilde{w}\,h\,d\bm{y}\,.$
@@
$\int\limits_{\Gamma}\tilde{w}\,h\,d\bm{y}=\int\limits_{\Gamma}\underbrace{w(\bm{y};\infty)}_{=1}\partial_{n}\tilde{w}\,d\bm{y}=0,$ (68)
@@
$=\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}\,d\bm{y}+\mu\int\limits_{\Gamma}\tilde{w}^{2}\,d\bm{y}.$ (69)
@@
$\frac{\eta_{1}(\eta_{1}-2\pi/\mu)}{|\Gamma_{1}|}\leq\frac{\eta_{2}(\eta_{2}-2\pi/\mu)}{|\Gamma_{2}|}\,$, (71)
@@ II.7 Large-reactivity limit
$\mu\biggl(\frac{1}{C(\mu)}-\frac{1}{C(\infty)}\biggr)=2\pi\sum\limits_{k=0}^{\infty}\frac{\mu[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}+\mu}\propto\ln(\mu)\to\infty,$ (81)
@@
$\times\underbrace{\sum\limits_{k=0}^{\infty}\Psi_{k}^{N}(\bm{y}_{1})\Psi_{k}^{N}(\bm{y}_{2})}_{=\delta(\bm{y}_{1}-\bm{y}_{2})}=\int\limits_{\Gamma}h^{2}(\bm{y})\,d\bm{y}=\infty.$
In other words, the squared harmonic measure density diverges logarithmically due to the edge singularity. This is exemplified by Eq. (75) that provides an explicit form of $h(\bm{y})$ for elliptic patches. In summary, even though the coefficients $\Psi_{k}^{N}(\infty)]^{2}$ determine the relative contributions of the Steklov eigenfunctions $\Psi_{k}^{N}$ to the reactive capacitance via Eq. (25), in analogy to the coefficients $F_{k}$ in Eq. (14), the latter are summed to $1$ that facilitates their interpretation as relative weights of the Steklov eigenfunctions $\Psi_{k}$ to the reactive capacitance $C(\mu)$.
@@ II.8 Bounds for the principal eigenvalue
$\int\limits_{\Gamma}\Psi_{0}\,d\bm{y}\leq\sqrt{2\pi c_{n+1}\mu_{0}^{n}}\,,\qquad n=1,2,\ldots$ (89)
@@
where $E(x)$ is the complete elliptic integral of the second kind. Substituting the values of $C(\infty)=2/\pi$, $c_{2}$ and $c_{3}$ into Eqs. (84, 85), we get three upper bounds: $1.263$ (Payne’s bound), $1.178$ ($n=1$), and $1.170$ ($n=2$), whereas the numerical value of the principal eigenvalue is $1.159$. One sees that the bound becomes tighter as $n$ increases, as expected from the spectral form (21). Using the numerical value $F_{0}\approx 0.978$, we get the lower bounds $1.152$ and $1.157$ for $n=1$ and $n=2$, and the last bound is very close to the numerical value of $\mu_{0}$. We recall that $c_{2}$ and $c_{3}$ for arbitrary patches are expressed via Eq. (18) in terms of the function $\omega_{\Gamma}(\bm{x})$ from Eq. (19). In Appendix B, we compute $\mathcal{A}_{\Gamma}$ for rectangular and elliptic patches; in the latter case, we also compare the upper bounds (84) and (87).
@@ III.1 Steklov eigenpairs
We complete this discussion by illustrating the difference between the Steklov problems (3) and (23). Figure 4 presents the first 9 eigenfunctions $\Psi_{k}^{N}|_{\Gamma}$ for the elliptic patch of semiaxes $1$ and $0.5$. Their comparison with panels of Fig. 3 indicates that, apart from the principal eigenfunctions $\Psi_{0}$ and $\Psi_{0}^{N}$, which are different, the other eigenfunctions look quite similar, up to matching their indices. Moreover, antisymmetric eigenfunctions are actually identical, up to their signs, e.g. $\Psi_{1}^{N}=-\Psi_{1}$, $\Psi_{3}^{N}=-\Psi_{2}$, etc. This is not surprising (see Appendix A for details): as the kernels $\mathcal{G}(\bm{y},\bm{y}^{\prime})$ and $\mathcal{G}^{N}(\bm{y},\bm{y}^{\prime})$ differ by a function $\omega_{\Gamma}(\bm{y})/|\Gamma|+\omega_{\Gamma}(\bm{y}^{\prime})/|\Gamma|-\mathcal{A}_{\Gamma}$, which is symmetric on the elliptic patch, the antisymmetric eigenfunctions of these kernels should be identical.
@@
Figure 7: First 15 Steklov eigenfunctions $\Psi_{k}$ restricted to the patch formed by two disks of radii $1$ and $0.5$, separated by distance $0.5$ (the associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the titles). The eigenfunctions were obtained by using a mesh with 964 triangles and 536 nodes.